Determine the equation of the line whose slope is -4 and the line crosses the y-axis at
(0, -2/3).
1
Expert's answer
2020-09-22T16:05:00-0400

Let the general equation of line be-

y = mx + C

Given : slope(m) = -4, so.

y = -4x + C

Since, the line passes through the point (0, -2/3), therefore, this point will satisfy the equation, i.e.,

-2/3 = -4(0) + C

Intercept(C) = -2/3

Equation of line : y = - 4x - 2/3